A rectangle has a width of 3X+5, the length is twice this. If the perimeter is 228 find the dimensions and area.
W ______, L __________ Area ___________


Answer by mananth(12270)

width of 3X+5,
the length is twice this
2(3x+5)
Perimeter = 2(L+W)
228 = 2((3x+5)+2(3x+5))
228=2(3x+5+6x+10)
228=2(9x+15)
228=18x+30
-30
198=18x
/18
x=11
W _38_____, L ___72_______ Area ____2736 sq. units_______